Lucido Products markets two computer games: Claimjumper and Makeover. A contribution format income statement for a recent monWhat is the overall contribution margin (CM) ratio for the company? 1% Overall CM ratioWhat is the companys overall break-even point in dollar sales? (Do not round intermediate calculations.) Overal break-even pPrepare a contribution format income statement at the companys break-even point that shows the appropriate levels of sales f
Answer #1

1.

Overall Contribution margin ratio = Overall Contribution margin/Total sales

= 132,750/177,000

= 75%

2.

Overall Break even point ($) = Fixed cost/Overall Contribution margin ratio

= 87,975/75%

= $117,300

3.

Sales ratio = 118,000:59,000

= 2:1

Contribution margin ratio = Contribution margin/Sales

Contribution margin ratio of Claimjumper = 81,600/118,000

= 69.15254237%

Contribution margin ratio of Makeover = 51,150/59,000

= 86.69491525%

Contribution format income statement

Claimjumper Makeover Total
Sales 117,300 x 2/3 = 78,200 117,300 x 1/3 = 39,100 117,300
Variable cost - 24,123 - 5,202 - 29,325
Contribution margin 54,077 33,898 87,975
Fixed cost - 87,975
Net operating income 0

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubts. Thanks.